OpenStudy (anonymous):

Solve for the variable in this equation x/2-2/3=1/4 Choose the answer that belongs x= 1 5/6 or x=11/12 or x=-5/6 or x=2

OpenStudy (anonymous):

x/2-2/3=1/4 x/2-2/3+2/3=1/4+2/3 x/2=11/12 (x/2)(2)=(11/12)(2) x=11/6

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):